OpenStudy (anonymous):

Which is the set of roots for the equation? x2 + 3x - 18 = 0 A. {18,-1} B. {-6,3} C. {-2,9} D. {9}

OpenStudy (mertsj):

Can you factor that polynomial?

OpenStudy (anonymous):

I am really stuggling in math all together so I can honestly say No Im not sure how to do that and that is even after reading and rereading the lesson

OpenStudy (mertsj):

(x+6)(x-3)=0

OpenStudy (mertsj):

If the product is 0 then one of the factors has to be zero so write x+6=0 or x-3=0 and solve each equation.

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

I am sorry for not being able to get it I am having such a bad rough da

OpenStudy (mertsj):

So can you solve the equation: \[x+6=0\]

OpenStudy (mertsj):

And how about this one? \[x-3=0\]

OpenStudy (anonymous):

No I am so lost

OpenStudy (mertsj):